OpenStudy (anonymous):

confirm the identity 2cot2x=cotx-tanx

OpenStudy (anonymous):

Put in terms of tan: = 2 / tan(2x) Double angle trig identity: = 2 / [2tan(x) / [1 - tan^2(x)]] Divide the fractions: = [2(1 - tan^2(x))] / 2tan(x) Divide numerator and denominator by 2: = (1 - tan^2(x)) / tan(x) Separate into two fractions: = 1/tan(x) - tan^2(x)/tan(x) Simplify: = cot(x) - tan(x)